in a random sample of 50 Lenawee county residents, the mean amount of property tax owed was $ 1,957. Assume that the population standard deviation for the amount of property tax was $256. Construct a 96% confidence interval for the mean amount of property tax paid by Lenawee County residents.
Answer #1

Solution :

Given that,

Sample size = n = 50

Z\alpha/2 = 2.054

Margin of error = E = Z\alpha/2* (\sigma /\sqrtn)

= 2.054 * ( 256/ \sqrt 50)

Margin of error = E = 74.36

At 96% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

1957 - 74.36 < \mu < 1957 + 74.36

1882.64 < \mu < 2031.36

(1882.64 , 2031.36)